OpenStudy (anonymous):

heather invested 20,000 in an account that pay 6% simple interest if she invested the money for 10 years how much will she have

OpenStudy (anonymous):

Interest will be 12000 so she will get 32000

OpenStudy (anonymous):

for 10 yeasr?

OpenStudy (anonymous):

Yes, the working is as follows P x T x R 20,000 x 10 x 6 1200000 Simple interest = ---------- = --------------- = ---------- = 12000 100 100 100 Amount = P + SI = 20000 + 12000 = 32000 Clear now ??? If yes, pls click the GOOD ANSWER button ;)